OpenStudy (anonymous):

could someone please help me wiith thois question im very confused,( find the slope of y=-5x+2)

OpenStudy (mle):

Hey there! and Welcome to OpenStudy! :) this is in "y-intercept form" which is when y = mx + b where m is the slope and b is the y intercept so y = mx + b y = -5x + 2 ^ m is the slope so can't you tell what the slope is now? :)

OpenStudy (anonymous):

yes i can thanks you for explaining to me :]

OpenStudy (mle):

Glad I could help :)

OpenStudy (anonymous):

? what should i substitute for x ?

OpenStudy (mle):

nothing, I don't thing you need to for this question :)

OpenStudy (anonymous):

so y = 7?

OpenStudy (mle):

are we on the same question?

OpenStudy (anonymous):

yes :[ i dont understand linear equations very well

OpenStudy (mle):

it's alright the slope would simply be the Negative 5

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

my answer choices arre 2,5,7 and 4

OpenStudy (anonymous):

so just positive 5

OpenStudy (mle):

hmm that's weird, I Don't think that's correct :-/

OpenStudy (anonymous):

oh i put a - in between 5 and = my bad it would be positive five right ;]

OpenStudy (mle):

oh, then yes :)

OpenStudy (anonymous):

lol thanks again

OpenStudy (anonymous):

sorry need help again 4 question later :/

OpenStudy (anonymous):

Which equation represents the graph shown?

OpenStudy (mle):

I can't seem to open the file?

OpenStudy (anonymous):

ok nvm then

OpenStudy (anonymous):

if the line is in bettween 1 and 2 would the y intercept be 1.5?

OpenStudy (anonymous):

and the x axis is in -3 in fourth quadrant

OpenStudy (mle):

I suppose so :) 

OpenStudy (anonymous):

hey man.. im no help but i HAD to say i love the picture :D

OpenStudy (anonymous):

ok thanks mLe and lol FlamePlays Mc

OpenStudy (anonymous):

how would you find the inequality in this equation

OpenStudy (anonymous):

Solve the inequality for b: -9b ≤ -36

OpenStudy (mle):

divide both sides by Negative 9 and flip the Sign from facing < To >

OpenStudy (anonymous):

oh ok that was really simple i feel dumb now

OpenStudy (anonymous):

wich one of these would it be ] b ≤ - 4 b ≥ 4 b ≤ 4 b ≥ - 4

OpenStudy (anonymous):

is it the top one?

OpenStudy (mle):

negative times negative is Positive and we flipped the sign so now it is facing >

OpenStudy (anonymous):

so it would be the second one :]

OpenStudy (mle):

mhmm :)